Cassini Significant Events 08/25/05 - 08/30/05

September 2, 2005

(Source: Cassini Project)


Thursday, August 25 (DOY 237):

Orbit trim maneuver #29 (OTM-29), the Titan 6 cleanup maneuver, was successfully performed today. The main engine burn began at 11:31 am PDT. A "quick look" immediately after the maneuver showed the burn duration was 9.3 seconds, giving a delta-V of 1.4 m/s. All subsystems reported nominal performance after the maneuver.

The Cassini Integrated Test Laboratory team is currently hosting a member of the Cosmic Dust Analyzer (CDA) instrument team at JPL. CDA has brought their engineering model to be connected to the ITL for flight software testing. The hardware will be here till mid-September.

Friday, August 26 (DOY 238):

The S15 sequence leads coordinated a test to be performed in the ITL of the Radar Titan 8 observations and special playbacks of data for that flyby. Analysis is underway to verify that the tests were successful.

Nine Instrument Expanded Block (IEB) files were uplinked to the spacecraft in support of S14. Sequence leads were able to verify that 8 of the 9 IEBs executed nominally and the readouts of telemetry were as predicted. There were no dropped packets for these files. Cassini Plasma Spectrometer (CAPS) team members will verify the ninth file. S14 will begin execution on Tuesday of next week.

Saturday, August 27 (DOY 239):

An additional IEB file and the S14 background sequence went up to the spacecraft today.

Monday, August 29 (DOY 241):

The 37th Cassini Project Science Group Meeting got underway this week at Imperial College, London, England.

The S17 Science Operations Plan (SOP) Update process officially began today. The SOP Update is a phase in the sequence development process where scientists and other operations team members are allowed to make limited observation and activity design changes from what was developed in SOP Implementation. This allows for late breaking discoveries and other information learned in prior sequences to be incorporated into future observations that have not completed their sequence development. S17 was originally archived back in April of 2003 so it has been over two years since scientists have looked at this product. The kick-off meeting, led by the Science Planning Team Lead, lays out the ground rules and schedule which all team members must abide by in order to incorporate their changes.

SOP Update for S16 continued this week with the completion of the Spacecraft Operations Office and Science Planning Team analysis of the S16 sequence merge. Output products were posted to the program file repository for review.

A kick-off meeting was held for the S14 DOY 248-250 Live Update process for an Inertial Vector Propagator update and Radio Science Live Movable Block. The current version of the schedule for this process has Navigation delivering the orbit determination (OD) solution at 1700 PDT today, with Science Planning (SP) releasing its epoch time shift and GEOEPOCHS file 2 hours later, and Radio Science (RSS) starting its analysis at that point and continuing for the next 8 hours. Navigation is aiming for an 1100 PDT OD solution release, so SP and RSS can get an earlier start. The Go/No-Go meeting is scheduled for tomorrow so the instruments and SP are expected to have their analysis ready by then. Update: It's a go!

Spacecraft Operations Office, Navigation, and Uplink Operations files for OTM-30 have been delivered and placed in the program file repository. The maneuver approval meeting will be held today at 5:00pm PDT. Uplink begins Tuesday at 07:05 AM PDT. On-station time is 11:45 AM for a 1:05 PM burn.

Tuesday, August 30 (DOY 242):

The S14 background sequence began execution today. The sequence will run for 39 days and will conclude on October 8. During S14 there will be two targeted flybys, two dust hazards to be avoided, eight OTMs, one live movable block, and three live IVP updates. Forty-three Deep Space Network (DSN) tracks will be used to downlink over 49.7 Gb of data.

A Design Delivery Review was held for Multi-Mission Image Processing Laboratory software version D34. This delivery contains many small updates to Imaging Science Subsystem & Visual and Infrared Mapping Spectrometer uplink and downlink software modules which correct product labels and eliminate operations workarounds. The delivery was accepted by all projects and MGSS. The software will go on-line on September 9.

The S17 Sponge Bit meeting was cancelled today due to a delay in the release of the DSN allocation file. The file is necessary for Cassini Science Planning to determine if there have been any changes to the amount of data that the DSN is capable of receiving. If they can receive more, then data volume that was held as margin is given to the teams for science acquisition. The allocation file will be delivered by the end of the week with the Sponge Bit meeting to follow.

Orbit trim maneuver #30 (OTM-30), the apoapsis maneuver preceding Titan-7, was successfully completed today. The main engine burn began at 13:05 am PDT. A "quick look" immediately after the maneuver showed the burn duration was 91.35 seconds, giving a delta-V of 14.3 m/s. All subsystems reported nominal performance after the maneuver.

A news release has been issued regarding the age of the "tiger stripes" discovered on Enceladus. Cracked features approximately 140 km long, spaced about 40 km apart and running roughly parallel to each other, act like vents spewing vapor and fine ice water particles that have become ice crystals. This crystallization process can help scientists pin down the age of the features. For the text of the full release and other materials, go to the Cassini Web Site at http://saturn.jpl.nasa.gov.

Wrap up:

The most recent spacecraft telemetry was acquired Tuesday, August 30, from the Goldstone tracking complex. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Check out the Cassini web site at http://saturn.jpl.nasa.gov for the latest press releases and images.